) One mole of a monoatomic ideal gas at initial pressure of 30 atm and 600 K undergoes a rapid adiabatic free expansion from a vessel to another 50 times larger in volume. Find the change in temperature and the increase in entropy.
Answer #1

For rapid free expansion,
Q = 0
W = 0

So,
∆U = - W + Q = 0
As Temperature depends on the Change in internal Energy,
This means, ∆T = 0

For Entropy,
∆S = Cv * ln(To/To) + nR*ln(V/Vo)
∆S = 0 + 1 * 8.314 * ln(50)
∆S = 32.5 J/K